(1 point) Let u= (3, 2) and v = (1,5). Then u +v=< 7 U-v=< -30 =< u. V = and || 0 ||
Answer #1

U 243, 27, V=<1,57 utva < (3+1), (2+ 5)2 (utv) = <4,7) U-V= < (3-4), (2-5)) una < 2,-3> - 3r- (-3,-15) u.rs $3637 265)} uov=